if two sides of a square are increased by five feet, the area of the field would be increased by 245 ft. find the area of the original square?

Answer by checkley77(12844)

x*x=x^2
(x+5)^2=x^2+245
x^2+10x+25=x^2+245
10x=245-25
10x=220
x=220/10
x=22 is the size of the original square.
Proof:
22^2=484
(22+5)^2=22^2+245
27^2=484+245
729=729